OpenStudy (anonymous):

factor the polynomial completely: 343+x^3

OpenStudy (amorfide):

are you sure this is typed correctly/the right question?

OpenStudy (anonymous):

yup o.o

OpenStudy (helder_edwin):

u have to use this general case \[ \large a^3+b^3=(a+b)(a^2-ab+b^2) \] for that u need to find z such that z^3=343

OpenStudy (anonymous):

its 7 right? o.o

OpenStudy (anonymous):

No, the question isn't factorable.

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

so \[ \large 343+x^3=7^3+x^3=? \]